NAOJ GW Elog Logbook 3.2

trying to have it better centered on the valve's window.
The height respect to the stack doesn't change. The lateral position is now 1.4 cm
west.

According to the records it should be a LISM mirror with the two faces flat.
So I reflect a green pointer on the mirror and collect the two reflections
from the two sides of the mirror. I let the beams propagates over about
15 m. I can't see any difference between the beams along the propagation.
So the two sides of the mirror are both flat (or both curved with a large
curvature). The distance between the two reflections is about 2.7 cm and does
not change in a significant way along the 15 m, so the mirror does not
have a large wedge.

I define as the tube axis the axis defined by the window on the valve between the NM2 chamber and the
tube and first target in the tube. So I placed an He-Ne beam on that axis.
I doubt to be able to defined that axis to better than 0.5 cm in position and 1 cm over ~10 m in angle.
This axis intercepts the stack in the NM1 22.2 cm in height and 1.8 cm to the west.
I estimate that I have about 1 cm error on this numbers.

In the second picture, there are totally five layers of the picomotor driver, except the first layer(command layer) is for pad, external port, which cannot connect to picomotors, the other four layers each has 3 ports, so now we have 12 ports for BS chamber and input mirror, and actually we still need one more port.
Yesterday, we test all the picomotors connected to this driver, all of them worked well with the command send from the computer, we tried with pad for control the five picomotors for input mirror, they worked well, but there is one thing we cannot understand is that on the pad(3rd picture), on the top of it, there are two axis, x and y, some of the picmotors work with x axis and others work with y, we cannot figure out what these axis means.
The other problem is that on the pad, the maximum layers we can control is 3, so if we have more than three layers, we cannot control with the pad.
For the PR chamber, we are going to use the other driver with 1 command layer with other 3 layers totally 9 ports to control, we will try to set a new IP address for the new driver.
Also we got a set of driver from Kamioka yesterday, one command layer with other 2 layers. So now we have enough command layer(one for BS and IM, one for PR, and one for EM), the port layer(5 for BS and IM, 3 for PR,1 for EM), we still lack of one port layer if all the driver we have now work well.


From the second picture can also see that now, since we only need to control five picomotors, so we only put two layers(except the first layer).
I list this simple command to move the picomotor for easy checking later: now for us there are only a1,a2 two layers, and each layer have 0,1,2 motor
chl--> to see which motor is working in each layer
chl a1=1--> change the working motor of a1 to 1
vel--> check the velocity of each motor
vel a1 1=1000 -->change the velocity of a1 motor 1 to 1000
rel a1 1000 g --> move a1(the motor which move is defined by the second command)
Next step we are going to make and change the all the cables of the BS and PR chamber, the BS chamber can use the same driver with the input mirror, just add more layers. But for the PR chamber, we need to use a new driver, which means we need to set up another IP address for that driver. There is also one driver left then for the end mirror. Also we need to check if the picomotors also work well with the pad.

I put the 500mm focal length lens after the lens we have now, shows in the first picture, under this circumstance, the beam is still circular after the bs mirror of the MZ, but after the mirror of the MZ, it becomes two separated circular beams because the second surface reflection of that mirror. These two beam goes along the path, at some point around the last mirror on the bench, they overlapped. So before the beam goes into the window, it is a really good beam, but after the window it’s separated into two again, so at first I thought maybe this is the reason of the distortion.
Then I blocked the reflection beam from the mirror on the bench, only left the main one, the second beam after the window is still there, so this second beam is produced by the reflection of the window.
Then inside the chamber, the beam after the window, only the main one arrives at the first dichoric mirror, the other one hit the frame of the suspension. So till here, we only have one circular beam, then I checked beam between the PR chamber and the BS chamber, it turns into ellipses. We talked with Matteo and we thought the problem is from the dichoric mirror which combine the green and the infrared. But since now we are not in a very good alignment, we will do the alignment again, and then check the shape of the beam. If it is still the same, we will change the dichoric mirror and to see if we can have some positive result.
For choosing the lens to reduce the beam, now I don't know what kind of focal length I should use. I think I will do one calculation with the measurement I did, got some result, and to see the roughly range of the focal length I can use, then put them there and do the measurement again to find better combination.
About re-align the beam, Matteo also gave us some steps to do that:
(All the things we need to move including, the turning mirror for green, the dichoric mirror for combining the green and infrared, the 2 inch mirror of the telescope, the PR suspension mirror, the BS suspension mirror. The mirrors are in the order of the beam pass through.)
1.Put the target in front of the 2 inch mirror, move the turning mirrors to center it.
2.Put the infrared card in front of the PR suspension mirror, center the green (and infrared?) with moving the 2 inch mirror.
3.Target in front of the BS suspension mirror, move the PR suspension mirror.
4.Check if the beam is centered at the 2 inch mirror(?)
5.Center the infrared beam on the BS suspension mirror with moving the dichoric mirror.
6.Check at the 2 inch mirror, the green and the infrared surpass or not.
Repeat these steps, until everywhere is centered.

Firstly, I tried the optical fiber used to send the PD signal,connected it to the camera at the end room, and connected it to the screen in the central room. There is only some flash when I turned on and off the receiver, but seems not capable to get the real signal from the camera.
Then we found out there is another system like a recorder that can be used. I plugged the recorder with the camera and the output to the screen at the end room. It works pretty well. And also there is a sticker with an IP address on it. We found the right internet cable for it, and tried to connected it with the computer in the center room. We successfully connected it, but cannot see the video signal.
Then we checked the old TAMA board which can convert the video into optical signal, send through the fiber and there is another this kind of board in the central room to convert it back, then connect to the screen, then we should see the video on the screen. At first, we tried with the cable 1-13 and 1-14, but we are not able to see anything from the screen. Then we tried to change the cable into 1-15 and 1-16, then we got the video in the central room.
The first picture is taken on the screen near the camera at 290m, the second picture is taken in the screen in the central room, both taken by my phone. It seems there are some loss during the signal transmission. Then we adjust the camera's position and also the focal lengths, and inject the red light into the arm, the third and the forth pictures are taken at 290m and central room.
The fifth picture is the board we are using now.

I upload some slides I have prepared to summarize the picomotor control issue we are facing.

The right dimension should be about 1mm when it arrives at the 2 inch mirror and a little bit larger like around 1.4mm at the FI, so the beam waist should almost at the mirror.
With these new conditions, I found some combination, chose the best one in the size and position, and check the position of the lens not overlap the mirror position we already mounted on the bench. The result shows as below.
The origin is set at the front surface of the BS mirror as I mentioned in my last entry.
L1: z=48.5cm f=500mm(already mounted)
L2: z=173.5cm f=300mm
L3: z=236cm f=350mm
With these lens, the final beam size at the FI is 1.368mm and the size of the beam waist is 0.97mm at the position of 648.57cm with the mirror's position at 648.15cm.
Also check again with JamMT, it shows the same result with the calculation.

We measured the infrared beam size last Friday and calculated the beam waist size and position:(set the front surface of the BS mirror as origin) the position of the beam waist is z=-12.0623 cm, size is 98.0277 um.
Then we measured the distance between the origin and FI in the PR chamber, and between the origin and 2inch mirror in the BS chamber.
The calculation process is as below:
The distance between the O and FI(z1)=277.5(The path length on the bench)+63.65(From the last hole of the bench to the window)+13.5(the depth of the viewport)+26.5(the distance between the chamber wall and the first mirror)+5.5(the length between and the FI)=386.65cm
The distance between the O and mirror(z2)=z1+16(the FI length plus the distance between FI output and the second mirror)+16(the second mirror to the dichoric mirror)+37(dichoric mirror to the chamber wall)+162(the distance between two chambers)+15(the BS chamber wall to the chamber bench)+57.5(the edge of the bench to the mirror)=648.15cm
The first lens is already put with focal length of 500mm at 48.5cm. Then there are other conditions need to be satisfied.
1.The beam waist should be around the 2inch mirror in BS chamber, and the size should be less than 103 um
2. The size of the beam at FI should also less than 103um.
3. All the lens should on the bench which means should located in the position z<=270 cm, because the last mirror is at 270cm.
I am trying to find a good combination with other two lens, but until now I did not get any result, I will try with more combination or I will add one more lens. Also today I measured the green beam path, but actually I did not know if I should trust the measure result or not, since the green beam is almost the same size or even larger than the size of the beam profiler, so I will try to put another lens first and try to measure it again.

Participants: Eleonora, Manuel, Yuefan, Raffaele
Today we have worked on the injection of the beam, exiting the telescope, in the south arm cavity.
1) We have moved the 2" mirror telscope in the BS chamber 5 cm closer to the PR in order to have a distance of 272.5 cm between the two telescope mirrors.
distance = f1 + f2 = 272.5 cm where f1 = 302.5 cm (see entry 419) and f2 = -30 cm (see entry 431)
2) We have used a spare TAMA PR mirror (ROC of about 9 km, almost flat) placed on a tripod, as steering mirror to send the beam in the corridor beside the pipe in order to check its features. We manage to send it into the end room and obseved its shape after 300 m. We remark that:
a) The beam is well collimated as it keeps roughly the same dimension after 300 m.
b) The beam moves a lot because of the air
3) We injected the green beam through the first viewport before NM2 chamber and tried to observe it on a target placed at 10 m, without success. We changed the beam direction moving the yaw and the pitch of the last telescope mirror.
4) We injected a red beam in the viewport and we were easily able to make it impinge on the target at 10 m. We observed that, in order to be centered both on the target and on the viewport the beam has to start from a position about 4 mm higher than the green beam position (21.3 cm from the stack top plate)
Conclusion: It is not clear the reason why we cannot see the green beam on the target at 10 m. The only two possible reasons we could imagine are that we were not able to properly align the beam or that it is not enough transmitted by the viewports or by the temporary mirror installed in the NM2 chamber (a former LISM mirror ). Effects of a possibile antireflective coatings for the infrared light on the green light should be investigated.
Attached pictures:
1- Green beam at 150 m
2- Green beam at 300 m
3- Red beam impinging to the target at 10 m

I have measured transfer functions and spectra for the PR mirror.
- Figure 1: TF when injecting noise in YAW
- Figure 2: TF when injecting noise in PITCH
- Figure 3: spectrum YAW (comparison open and closed loop)
- Figure 4: spectrum PITCH (comparison open and closed loop)
NB. It seems that there is some acoustic noise: I did the measurement with the light switched off but the chamber was open and I didn't put any boxex around the laser and the PSD
While looking at the error signals in the past days I have noticed that often some strong oscillations at 1 Hz and 3 Hz were present (see figure 5). We observed it also in the past and we were suspecting the stack vibratiion introduced by the steering mirrors used for the optical lever ( see entry 276: the attached spectra taken last july are very similar to those of fig 5)
Since this oscillation is not present when the intermadiate mass is well centered on the magnet older plate (as it was just after the mirror installation or today after that I performed the check on the yaw motion) and it appeared after we tried to move as much as possible the yaw picomotor yesterday, I think it is caused by the intermediate mass touching the magnet holder plate.

In order to better quantify the yaw motion of the PR suspended mirror I made the optical lever reflected beam imping on a ruler put at a distance of about 88 cm from the mirror.(See attached picture 1)
I started from one side of the picomotor range (left column of the second attached picture) and I kept turing the screw as far as I was able to see a beam movements on the ruler.
I had to stop well before that the end of the range of the picomotor (right column of the second attached picture) since the beam was not moving anymore (and the mirror was oscillating a lot)
Even if it difficult to clearly see it, I think that the intermiate mass was touching the magnet holder plate preventing a further motion.
The total motion of the beam on the ruler is about 1.5 cm which corresponds to 8.5e-3 rad.

After the problem reported yesterday here. I have replaced the picomotors GPIB driver with a similar one that I have found in the south end room.
At the beginning it seems to work fine and I was able to move the 3 degree of freedom of the 2'' telescope mirror in both direction both with the pad that using the PC.
Then, I have noticed that while connecting more than one cable to the driver output ports, it starts showing some problems.
Its behaviour becomes very strange and it is no loger possibile to move properly the selected channel neither from remote nor with the pad.
For exemple: if I send a remote command to move a specific channel the driver display shows correctly the name of the selected channel but it cannot move it and will instead make other picomotor ( connected to a different channel) to produce some noise.
I could not find any repetable pattern in such kind of strange behaviour while trying to test different combination of slots and cables.
I only noticed that everything is working fine when only one cable is connected.
Since Raffaele suggested that it can be a cabling problem, I spent some times to check the connections but I haven't find anything wrong so far.
In the attachement there is a picture of the connection we made between the 2" telescope picomotor and the invacuum cable.

Participants: Matteo Tacca, Eleonora
Today we installed the 2" telescope mirror in the BS chamber. The infrared and green beams have been superposed. They impinge the first telescope mirror and are reflected on the second one (PR suspended mirror). Beams are both hitting the PR mirror almost in the center. We observed that the green beam is slightly elliptical (also before hitting the first telescope mirror)
Then we tried to send the beams back on the BS mirror but we didn't manage to move the yaw of the PR mirror enough to do that.
We spent some time investigating this issue.
Some remarks:
1) The PR mirror is very sensitive to pich motion: few turns of the pitch picomotors screw are enough to remarkably change the vertical angle of the beam.
2) The yaw picomotor screw works fine as we can observe the relative slide to move (in both directions) when turning the screw.
3) We checked that the mirror is free to move and it not touching the earthquake stops
4) It was possible to observe the pitch motion using the error signal of the optical lever. The displacement is big enough to make it sweep the all the signal range (from -5V to +5V). Assuming a calibration of 2.2e-4 rad/V this means that the yaw motion achievable is al least larger than 2.2e-3 rad.
5) Another critical point is the intermidiate mass. During the mirror istallation we noticed that one side of it was almost touching the magnet holder plate. We cannot rule out the possibility that the yaw motion is limited by this.

Participants: Matteo Tacca, Eleonora
Today we moved the BS suspension in order to have the reflecting surface of the mirror 10 cm south with respect to the center of the chamber, according to the telescope design.
We have also installed the fixed mirror support and connected the in-vacuum and out-of-vacuum cable to move the picomotors. We had some issues in moving the pitch and the translation picomotors in one direction (the other direction works fine).
The picomotors seems to receve signals from the drivers since they make a noise but most of the time they don't move and sometimes they move very slowly.
NB: Ony few days ago I I have checked that the 3 picomotors can work in both directions (entry 420)
We made many checks in order to find out the cause problem:
1) We checked carefully all the connections in and out of the vacuum chamber. They were all fine.
2) We changed the picomotor with a spare one but it didn't solved the problem. So it is not a problem of the picomotor itself.
3) We have tried to control them by PC using labview: nothing changed.
4) We have changed the pad of the driver: nothing changed.
5) We have changed the slot at which the cable was connected in the driver: nothing changed.
6) We have inverted the pins of pitch and yaw in the in vacuum connector and this time the pitch was working fine in both directions while the yaw was now showing the issue.
This last test made as to suspect the driver which, by the way, showed some other problems ( i.e the knob to change the controlled slot gets often stuck and it is necessary to switch it on and off).
I found a similar driver in the instrument shelf at the first flor but it didn't switch on.
Tomorrow we will:
1) Check the driver on other picomors, as those of the PR suspension which were also succesfully tested before.
2) Look for other GPIB picomotor drivers in TAMA. (NB according to their label these drivers have been bought in 1999)

Today I restored the optical lever of the telescope suspended mirror in PR chamber. Error signals look fine even if, since the chamber is open, they are a bit noisy than usual due to the light.
In order to close the loops I had to increase a bit the gain of the yaw loop and change both the loop signs. I have also tweaked the driving matrix to reduce the exicitation of yaw when injecting noise on pitch. In the previous configuration I have added a rotation of the sensing of 0.04 rad (as reported here), now I have removed it since it seems to be not needed anymore.
I will upload soon spectra and transfer functions.

Today while tiding up the area close to the BS chamber I noticed that there was same air leaking from a pipe between BS and NM2 chamber (see attached picture).
As soon as I touched the pipe tying to tight it better, it came off, loosing a lot of air. With the help of Matteo, Fujii and Ishizaki-san we managed to reconnect and tight it. Now the pipe seems well attached but there is still some leakege even if it is smaller than it was before.
It has to be checked by an expert as soon as possible.

Then install the reflect mirror just after the dichoric mirror temporarily to adjust the reflect beam of infrared to goes the position it should be at the optical bench as the new optical scheme.
We also check the power loss of the beam goes through the Faraday and comes back to the optical bench, it is about 6% and 10%.
The we put back the middle part of the PR chamber, since then the beam will first go through the thick window mirror, if it is not exactly perpendicular to the mirror, there is at least some shift of the beam. And also it seems the mirror also is a wedge mirror, now the beam in the chamber moves a lot. One of the idea is to move the last mirror before the beam goes into the chamber on the bench, but then there is the risk that the beam will be move out from the mirror. Matteo will work on this in these days.
And on Friday, we plan to measure the green beam, and do some calculation to decided the lens we are to use, and we should buy them, since now in the lab the shortest focus length is 100mm.
trying to have it better centered on the valve's window.
The height respect to the stack doesn't change. The lateral position is now 1.4 cm
west.
the He-Ne laser beam goes through the input mirror. Also in this case I centered the He-Ne laser beam
on the the first target in the tube and on the window at the entrance of the tube.
In order to do so the injection point from the back of the suspension is 21.8 cm in height and 2.4 cm
to the west.